# JANET: Joint Adaptive predictioN-region Estimation for Time-series

### v1

@@ 4.3 Validity of ICP with permutations
In Theorem 1 we establish that under mild assumptions on ergodicity and small prediction errors, as defined in A.2.1, we can achieve approximate validity in the case of a single time series whilst using the permutation scheme from Chernozhukov et al. (2018). In the case of IID data, we have exact validity guarantees shown in Theorem 2 and this exact validity can be applicable in the case of multiple independent time series.
Definition 1 (Randomised $p$-value)
We define the randomised $p$-value as:
$\hat{p}=\hat{p}(y):=\frac{1}{d}\sum_{j=1}^{d}\mathbf{1}\left(\mathcal{A}\left(Z_{tr},Z_{cal}^{\pi_{j}}\right)\geq\mathcal{A}\left(Z_{tr},Z_{cal}\right)\right)$
@@
where $\pi_{j}$ is the $j^{th}$ permutation of $\Pi$. We can invert the quantile to gain a prediction interval as in Eq. 1. Theorem 1 is adapted from Chernozhukov et al. (2018) to the inductive conformal setting.
@@
The proof can be found in Section A.2.
@@
Chernozhukov et al. (2018) demonstrated that conformity scores obtained via the permutations (Figure 1) offer a valid approximation to the true conformity score distribution for strongly mixing time series. Notably, stationary processes like Harris-recurrent Markov chains and autoregressive moving average (ARMA) models exhibit strong mixing properties Athreya and Pantula (1986b, a). Statistical tests designed for ARMA models, such as the Ljung-Box or KPSS tests, can be employed to assess the presence of strongly mixing. Our empirical findings suggest that even when stationarity is violated, approximate coverage can still be achieved (see Table 1).
@@ 5 JANET: Joint Adaptive predictioN-region Estimation for Time-series
A joint prediction region (JPR) typically controls the probability of the prediction region containing the entire true prediction sequence at a specified significance level $\epsilon$. This can be interpreted as the probability of observing at least one element of the true sequence outside the region, also known as the familywise error rate (FWER), namely:
[…]
When $K=1$, the $K$-FWER reduces to the standard FWER. By allowing a tolerance for errors, larger values of $K$ yield smaller, more informative prediction regions, which can be beneficial in decision-making scenarios where some degree of error is acceptable.
Remark: JPRs can be constructed in various forms. Whilst hyperspherical construction is a common choice, it may not facilitate reasoning about individual time steps in the horizon. Although it is possible to project the hypersphere onto a hyperrectangle to enable component-wise analysis, this results in a larger region and a loss of predictive efficiency Diquigiovanni et al. (2021); Wolf and Wunderli (2013).
We introduce JANET (Joint Adaptive predictioN-region Estimation for Time-series) to control $K$-FWER in multi-step time series prediction.
We adopt the notation for the multiple time series setting and further assume that the time series are univariate (i.e. $p=1$). In the case of a single time series, we use the same non-conformity scores but treat the permutations of the single time series as distinct, exchangeable time series. We propose two non-conformity measures that are extensions of a locally adaptive non-conformity score from Lei et al. (2018)
@@
for $i\in\{1,\dots,n_{cal}\}$ where, $K\text{-}\max(\vec{x})$ is the $K^{th}$ largest element of a sequence $\vec{x}$, $y_{i,j}$ is the $j^{th}$ entry of the target $Y_{i}$ and $\hat{y}_{i,j}$ is the $j^{th}$ entry of $\hat{f}(X_{i})$. Note that the prediction steps can be generated by any model $f$ (AR models with $H=1$ or multi-output models with $H>1$). The difference between Eqs. 5 and 6 is only in the scaling factors (denominators). In Eq. 5 the scaling factors $\hat{\sigma}_{1},\dots,\hat{\sigma}_{H}$ are standard deviations of the error, that are computed on the proper training sequence. These scaling factors account for differing levels of variability and magnitude of errors across the prediction horizon but do not depend on the history. Meanwhile, for Eq. 6 the scaling factors are conditional on the relevant history $X$.
@@
Note: In the multivariate case (i.e., $z\in\mathbb{R}^{p},\;p>1$), the entries of Eqs. 5 and 6, whose entries are $\left[\frac{|y_{i,j}-\hat{y}_{i,j}|}{\hat{\sigma}_{j}(X_{i})}\right]_{j}$ for $j\in\{1,\dots,p\}$. We then take the $-K\text{-}\max$, across all $p$ dimensions, and all $H$ steps in the time horizon.
@@
where $X^{*}$ is the sequence of lagged values (i.e. the history) for the unknown $Y^{*}$ that is to be predicted. The $-K\text{-}\max$ operation maps the multidimensional residuals to a singular score which allows for quantile inversion as done in Eq. 1. Further, by taking the $-K\text{-}\max$, each $\alpha^{K}_{i}$ (or $R^{K}_{i}$) is the value to rescale each $\hat{\sigma}_{j}$ by to provide intervals that cover all but $K$ of the predicted time points for a specific trajectory $Z_{i}\in\mathbf{Z}_{cal}$.
@@ 6 Experiments
We demonstrate the utility of our method, JANET, on single time series and multiple time series. For the single time series, we compare against Bonferroni and Scheffé JPRs. Further, we compare against bootstrapping methods. Note that for Bonferroni and Scheffé we can only train the models for complete coverage ($1$-FWE), whilst for Bootstrap-JPR we can vary $K$ (as with our method). Despite matching this level of flexibility in the choice of $K$, Bootstrap-JPR is much more computationally intensive than JANET. For the multiple time series, we compare against baselines (CopulaCPTS Sun and Yu (2024), CF-RNN Stankeviciute et al. (2021), MC-Dropout Gal and Ghahramani (2016)) that make stronger assumptions (different series being independent) than us on synthetic datasets as well as real-world data. Our generalised ICP method can lead to greater data efficiency whilst creating approximately valid prediction sets should independence be violated.
@@ 6.1 Single time series experiments
  • Bonferroni Correction (Dunn, 1961; Haynes, 2013): This approach controls the FWER, but it is conservative. We use bootstrapping-based methods to find the standard deviation of the prediction errors before applying the correction.
  • Scheffé-JPR (Staszewska-Bystrova, 2011): This statistical method assumes normality of errors but may not hold for prediction intervals.
  • Bootstrap-JPR (Wolf and Wunderli, 2013; English, 2024): This method is based on bootstrapping and lacks finite-sample guarantees and can be computationally demanding.
We use an ARIMA model as the learner for all methods in this section. Due to computational constraints over numerous simulations, we do not use neural networks for the main coverage experiments. Both Bootstrap-JPR and our proposed method, JANET, control the $K$-FWER (Lehmann and Romano, 2005), so we compare for different tolerance levels, $K$. We want to point out that Bootstrap-JPR can be conformalised in a naive way. Whilst conformalised bootstrapping can address finite-sample issues, it does not reduce computational cost. In contrast, JANET requires training the model only once per simulation. We compare both of our variants: JANET and JANET*.
@@ 6.1.1 Monte Carlo simulations
We generate data from an $AR(2)$ process with $\rho\in\{1.25,-0.75\}$ and evaluate empirical coverage across 1000 simulations. We compute the JPRs for $K\in\{1,2,3\}$, $H\in\{6,12,18,24\}$ and significance level, $\epsilon\in\{0.7,0.8,0.9\}$. For methods that cannot control the $K$-FWER, the corresponding results refer to $K=1$. The interval width of one JPR is calculated as the geometric mean of the widths over the horizon $H$. The average over all simulations per setting is reported in Tables 6, 7 and 8.
Tables 3, 4 and 5 present coverage results for $\epsilon$ and varying tolerances $K$, whilst Figures 4, 5 and 6 display coverage errors as bar plots. As expected, Bonferroni is conservative, particularly at larger $\epsilon$. Scheffé-JPR undercovers substantially, likely due to the normality assumption. Bootstrap-JPR and JANET* perform comparably. JANET often has smaller coverage errors but slightly larger average widths, as it treats errors uniformly across the history covariates. Additionally, Tables 6, 7 and 8 present empirical width results for various significance levels $\epsilon$ and varying tolerances $K$ and Figures 9, 8 and 7 plot empirical widths against forecast horizon ($H$) for various significance levels $\epsilon$ and $K=1$. Figure 3 compares coverage errors against widths, with the ideal method being close to zero error with minimal width within each colour group (representing the same tolerance $K$). Note that within a level of $K$, the points tend to cluster together. The left plot ($K=1$) includes more points per cluster as Bonferroni and Scheffé-JPR control for this tolerance level only. The analysis also shows that JANET* generally achieves good coverage with minimal width.
@@
We evaluate JANET on the U.S. real gross domestic product (GDP) dataset (U.S. Bureau of Economic Analysis, 1947). To address non-stationarity, we log-transform and de-trend the data as a preprocessing step. The resulting series is shown in Figure 10. Our task is to forecast the next $H=4$ quarters (equivalent to one year) and construct a JPR for the true sequence. We set the significance level to $\epsilon=0.2$. Due to the limited availability of real-world data, we create windowed datasets to increase the number of datasets for evaluating coverage. Each window consists of a sequence of 48 quarters (12 years) for JPR computation, followed by a true sequence of 4 quarters (1 year) for coverage assessment. It should be noted that this method for computing empirical coverage has two deficiencies, (i) there are only 100 series (created through windowing); (ii) the series are not independent of each other. Nevertheless, it still provides an assessment of the out-of-sample performance of the method. Table 1 shows the coverage results of the data. Similar to the experiment in Monte Carlo simulations in the previous section, the out-of-sample coverages for JANET and Bootstrap-JPR are close to the desired level at all tolerances $K$, whereas Bonferroni overcovers and Scheffé-JPR undercovers.
@@ 6.2 Multiple independent time series experiments
In this section, we focus on the setting with multiple independent time series as in (Stankeviciute et al., 2021; Sun and Yu, 2024). As previously noted, this scenario allows us to achieve exact validity guarantees, rather than approximate validity, due to the independence assumption. We compare JANET against the following methods:
  • CF-RNN (Stankeviciute et al., 2021) is designed for multi-output neural networks–the entire predictive sequence is outputted at once. It assumes conditional IID prediction errors, which may not always be accurate, especially if the trained model has not captured the underlying trend. Further, it relies on the Bonferroni correction (Haynes, 2013), which tends to be conservative.
  • MC-Dropout (Gal and Ghahramani, 2016) is a Bayesian method for neural networks that can provide prediction intervals that are often overly narrow intervals with poor coverage, indicating overconfidence.
  • CopulaCPTS (Sun and Yu, 2024) uses copulas to adjust for dependencies. However, it is data-inefficient since it requires two calibration sets and unlike JANET, it cannot adapt its regions based on history or handle different values of $K$ in the $K$-FWER control problem.
Following the evaluation approach used in the baseline methods, we report the "frequency of coverage" on the test set, which should not be interpreted as frequentist coverage, as the latter requires asymptotic repetition Vovk et al. (2023).
@@ 6.2.1 Particle simulation datasets
We evaluate our model on two synthetic datasets from (Kipf et al., 2018) using the same experimental settings as in (Sun and Yu, 2024). In both cases, we predict $H=24$ steps into the future based on a history of length $T=35$. Each time step is in $\mathbb{R}^{2}$. In the two setups, we add in mean-zero Gaussian noise with $\sigma=0.05$ and $\sigma=0.01$, respectively. We used two different predictors for the experiments. See Section B.2 for predictor model and training details. Table 2 shows results for the particle5 experiment $\sigma=0.05$. Under the EncDec predictor, our coverage is closer to the desired level. For RNN architectures, we observe slight under-coverage for JANET and over-coverage for CF-RNN, whilst other methods, especially MC-dropout, exhibit severe under-coverage. In the particle1 experiment (Table 2, $\sigma=0.01$), JANET achieves better coverage under both predictors. Other methods again show significant under-coverage, particularly MC dropout.

### v2

@@ 3.3 Validity of ICP with permutations
Theorem 1 states that under mild assumptions on ergodicity and small prediction errors, as defined in A.2.1, we can achieve approximate validity in the case of a single time series whilst using the permutation scheme from [10]. In the case of IID data, we have exact validity guarantees shown in Theorem 2 and this exact validity can be applicable in the case of multiple independent time series.
@@
where $\pi_{j}$ is the $j^{th}$ permutation of $\Pi$. We can invert the quantile to gain a prediction interval as in Eq. 1. Theorem 1 is adapted from [10] to the inductive conformal setting.
Theorem 1 (Approximate General Validity of Inductive Conformal Inference).
@@
Here, the terms $\delta_{1d}$ and $\gamma_{1d}$ correspond to the approximation error and failure probability in the empirical distribution of permutation-based conformity scores under weak ergodicity assumptions. The terms $\delta_{2m}$ and $\gamma_{2m}$ reflect the estimation error and failure probability when approximating the oracle non-conformity score using finite training data. Together, these bounds quantify the deviation from exact validity, and all vanish as $d,m\to\infty$. Full assumptions, definitions and the proof are provided in Section A.2.
@@
[10] demonstrated that conformity scores obtained via the permutations (Figure 1) offer a valid approximation to the true conformity score distribution for strongly mixing time series. Notably, stationary processes like Harris-recurrent Markov chains and autoregressive moving average (ARMA) models exhibit strong mixing properties [6, 5]. Statistical tests designed for ARMA models, such as the Ljung-Box or KPSS tests, can be employed to assess the presence of strongly mixing. Our empirical findings suggest that even when stationarity is violated, approximate coverage can still be achieved (see Table 1).
@@ 4 JANET: Joint Adaptive predictioN-region Estimation for Time-series
A joint prediction region (JPR) aims to capture the entire multi-step prediction sequence $Y=(y_{T+1},\dots,y_{T+H})$ with high probability. Traditionally, this is formalised by controlling the familywise error rate (FWER), which bounds the probability that any ($h$) of the $H$ components fall outside their respective prediction intervals ($C_{T+h}$):
$\text{FWER}:=\Pr\left(\exists\,h\in\{1,\dots,H\}\text{ such that }y_{T+h}\notin C_{T+h}\right).$
Equivalently, FWER corresponds to a joint coverage probability of at least $1-\epsilon$, where $\epsilon$ is the user-specified significance level. However, as the prediction horizon $H$ increases, FWER control becomes increasingly conservative, often producing excessively large prediction regions that reduce practical utility.
$K\text{-FWER}:=\Pr\left(\left|\{h:y_{y+h}\notin C_{T+h}\}\right|\geq K\right).$
When $K=1$, the definition recovers standard FWER. By allowing controlled violations, K-FWER enables tighter prediction regions that are more informative and practically actionable. For example, when $H=24$, allowing up to $K-1=2$ errors may be acceptable in high-uncertainty applications like macroeconomic forecasting. As such, K-FWER offers a more flexible and tunable approach to uncertainty quantification, especially in domains where small deviations can be tolerated in exchange for sharper predictions.
Remark. Joint prediction regions (JPRs) can be constructed in various geometric forms. A common approach is to use a hyperspherical region in $\mathbb{R}^{H}$, which may offer statistical or computational advantages. However, such regions are not decomposable across time steps, making it difficult to assess whether specific components of a forecast sequence lie within the region.
In contrast, we adopt a rectangular construction, where the joint region at $\epsilon$, the significance level (or $K-$FWER), $C_{1-\epsilon}=\prod_{h=1}^{H}C_{T+h,\epsilon}$ consists of marginal intervals along each forecast horizon. This structure offers a key interpretability advantage: one can track the realised sequence in real time and immediately identify whether a prediction lies outside the region at any step. For instance, if the forecast for day $T+h$ falls outside $C_{T+h}$, a decision-maker can already reason that the region may be invalidated, especially if $K$ or more such violations occur. This enables sequential, component-wise decision-making — a property not afforded by hyperspherical regions.
While it is possible to project a hypersphere onto a hyperrectangle to enable component-wise analysis, this leads to a looser region and reduced predictive efficiency [12, 42]. As such, rectangular regions are better suited to applications requiring transparent, step-level risk assessment.
Now, we introduce JANET (Joint Adaptive predictioN-region Estimation for Time-series) to control $K$-FWER in multi-step time series prediction.
We adopt the notation for the multiple time series setting and further assume that the time series are univariate (i.e. $p=1$). In the case of a single time series, we use the same non-conformity scores but treat the permutations of the single time series as distinct, exchangeable time series. We propose two non-conformity measures that are extensions of a locally adaptive non-conformity score from [27]
@@
for $i\in\{1,\dots,n_{cal}\}$ where, $K\text{-}\max(\vec{x})$ is the $K^{th}$ largest element of a sequence $\vec{x}$, $y_{i,j}$ is the $j^{th}$ entry of the target $Y_{i}$ and $\hat{y}_{i,j}$ is the $j^{th}$ entry of $\hat{f}(X_{i})$. Note that the prediction steps can be generated by any model $\hat{f}$ (AR models with $H=1$ or multi-output models with $H>1$). The difference between Eqs. 5 and 6 is only in the scaling factors (denominators). In Eq. 5 the scaling factors $\hat{\sigma}_{1},\dots,\hat{\sigma}_{H}$ are standard deviations of the error, that are computed on the proper training sequence. These scaling factors account for differing levels of variability and magnitude of errors across the prediction horizon but do not depend on the history. Meanwhile, for Eq. 6 the scaling factors are conditional on the relevant history $X$.
@@
Note: In the multivariate case (that is, $z\in\mathbb{R}^{p},\;p>1$), the entries of Eqs. 5 and 6, whose entries are $\left[\frac{|y_{i,j}-\hat{y}_{i,j}|}{\hat{\sigma}_{j}(X_{i})}\right]_{j}$ for $j\in\{1,\dots,p\}$. We then take the $-K\text{-}\max$, across all $p$ dimensions, and all $H$ steps in the time horizon.
@@
where $X^{*}$ is the sequence of lagged values (i.e. the history) for the unknown $Y^{*}$ that is to be predicted. The operation $-K\text{-}\max$ maps the multidimensional residuals to a singular score that allows quantile inversion as done in Eq. 1. Further, by taking the $-K\text{-}\max$, each $\alpha^{K}_{i}$ (or $R^{K}_{i}$) is the value to rescale each $\hat{\sigma}_{j}$ to provide intervals that cover all but $K$ of the predicted time points for a specific trajectory $Z_{i}\in\mathbf{Z}_{cal}$.
@@ 5 Experiments
We demonstrate the utility of our method, JANET, on single time series and multiple time series. For the single time series, we compare against Bonferroni and Scheffé JPRs. Further, we compare against bootstrapping methods. Note that for Bonferroni and Scheffé we can only train the models for complete coverage ($1$-FWE), whilst for Bootstrap-JPR we can vary $K$ (as with our method). Despite matching this level of flexibility in the choice of $K$, Bootstrap-JPR is much more computationally intensive than JANET. For the multiple time series, we compare against baselines (CopulaCPTS [38], CF-RNN [35], CAFHT [46], CCN-JPR [17], and MC-Dropout [19]) that make stronger assumptions (different series being independent) than us on synthetic datasets as well as real-world data. Our generalised ICP method can lead to greater data efficiency whilst creating approximately valid prediction sets should independence be violated.
@@ 5.1 Single time series experiments
  • Bonferroni Correction [14, 23]: This approach controls the FWER, but it is conservative. We use bootstrapping-based methods to find the standard deviation of the prediction errors before applying the correction.
  • Scheffé-JPR [37]: This statistical method assumes normality of errors but may not hold for prediction intervals.
  • Bootstrap-JPR [42, 16]: This method is based on bootstrapping and lacks finite-sample guarantees and can be computationally demanding.
We use an ARIMA model as the learner for all methods in this section. Due to computational constraints over numerous simulations, we do not use neural networks for the main coverage experiments. Both Bootstrap-JPR and our proposed method, JANET, control the $K$-FWER [26], so we compare for different tolerance levels, $K$. We want to point out that Bootstrap-JPR can be conformalised in a naive way. Whilst conformalised bootstrapping can address finite-sample issues, it does not reduce computational cost. In contrast, JANET requires training the model only once per simulation. See Appendix B for a detailed analysis of the computation costs. We compare both of our variants: JANET and JANET*.
@@ 5.1.1 Monte Carlo simulations
We generate data from an $AR(2)$ process with $\rho\in\{1.25,-0.75\}$ and evaluate empirical coverage across 1000 simulations. We compute the JPRs for $K\in\{1,2,3\}$, $H\in\{6,12,18,24\}$ and significance level, $\epsilon\in\{0.7,0.8,0.9\}$. For methods that cannot control the $K$-FWER, the corresponding results refer to $K=1$. The interval width of one JPR is calculated as the geometric mean of the widths over the horizon $H$. The average over all simulations per setting is reported in Tables 11, 12 and 13.
Tables 8, 9 and 10 in the Appendix present coverage results for $\epsilon$ and varying tolerances $K$, whilst Figures 5, 6 and 7 in the Appendix display coverage errors as bar plots. As expected, Bonferroni is conservative, particularly at larger $\epsilon$. Scheffé-JPR undercovers substantially, likely due to the normality assumption. Bootstrap-JPR and JANET* perform comparably. JANET often has smaller coverage errors but slightly larger average widths, as it treats errors uniformly across the history covariates. Additionally, Tables 11, 12 and 13 in the Appendix present empirical width results for various significance levels $\epsilon$ and varying tolerances $K$ and Figures 10, 9 and 8 in the Appendix plot empirical widths against forecast horizon ($H$) for various significance levels $\epsilon$ and $K=1$. Figure 2 compares coverage errors against widths, with the ideal method being close to zero error with minimal width within each colour group (representing the same tolerance $K$). Note that within a level of $K$, the points tend to cluster together. (refer to Figure 4 in the Appendix for all different $K$ and all significance levels $\alpha$). The left plot ($K=1$) includes more points per cluster as Bonferroni and Scheffé-JPR control for this tolerance level only. The analysis also shows that JANET* generally achieves good coverage with minimal width.
@@
We evaluate JANET on the U.S. real gross domestic product (GDP) dataset [40]. To address non-stationarity, we follow the preprocessing procedure used in the Bootstrap-JPR [42]: a log transformation followed by first-order differencing to remove trends. The resulting log-growth series is shown in Figure 3.
Our task is to forecast the next $H=4$ quarters (equivalent to one year) and construct a JPR for the true sequence. We set the significance level to $\epsilon=0.2$. Due to the limited availability of real-world data, we create windowed datasets to increase the number of datasets for evaluating coverage. Each window consists of a sequence of 48 quarters (12 years) for JPR computation, followed by a true sequence of 4 quarters (1 year) for coverage assessment. It should be noted that this method for computing empirical coverage has two deficiencies, (i) there are only 100 series (created through windowing); (ii) the series are not independent of each other. Nevertheless, it still provides an assessment of the out-of-sample performance of the method. Table 1 shows the coverage results of the data. Similar to the experiment in Monte Carlo simulations in the previous section, the out-of-sample coverages for JANET and Bootstrap-JPR are close to the desired level at all tolerances $K$, whereas Bonferroni overcovers and Scheffé-JPR undercovers.
@@ 5.2 Multiple independent time series experiments
In this section, we focus on the setting with multiple independent time series as in [35, 38]. As previously noted, this scenario allows us to achieve exact validity guarantees, rather than approximate validity, due to the independence assumption. We compare JANET against the following methods:
  • CF-RNN [35] is designed for multi-output neural networks–the entire predictive sequence is outputted at once. It assumes conditional IID prediction errors, which may not always be accurate, especially if the trained model has not captured the underlying trend. Further, it relies on the Bonferroni correction [23], which tends to be conservative.
  • MC-Dropout [19] is a Bayesian method for neural networks that can provide prediction intervals that are often overly narrow intervals with poor coverage, indicating overconfidence.
  • CopulaCPTS [38] uses copulas to adjust for dependencies. However, it is data-inefficient since it requires two calibration sets and unlike JANET, it cannot adapt its regions based on history or handle different values of $K$ in the $K$-FWER control problem.
  • CAFHT [46] uses ACI [9] for primarily heterogenous trajectories.. Just like CopulaCPTS, it is also data-inefficient due to additional tuning parameters and cannot adapt its regions based on history or handle different values of $K$ in the $K$-FWER control problem.
  • CCN-JPR [17] employ Normalising Flows to prediction regions that do not have any geometrical assumptions. Their appraoch restricts to using generative models with density estimation capabilities as predictors and can not be used with any other model as in our case.
Following the evaluation approach used in the baseline methods, we report the coverage on a hold-out test set.
@@ 5.2.1 Particle simulation datasets
We evaluate our model on two synthetic datasets from [25] using the same experimental settings as in [38]. In both cases, we predict $H=24$ steps into the future based on a history of length $T=35$. Each time step is multivariate, i.e., $\mathbb{R}^{2}$. In the two setups, we add in mean-zero Gaussian noise with $\sigma=0.05$ and $\sigma=0.01$, respectively. We used two different predictors for the experiments. See Section D.2 for predictor model and training details. Table 2 shows results for the particle5 experiment $\sigma=0.05$. Under the EncDec predictor, our coverage is closer to the desired level. For RNN architectures, we observe slight under-coverage for JANET and over-coverage for CF-RNN, whilst other methods, especially MC-dropout, exhibit severe under-coverage. In the particle1 experiment (Table 2, $\sigma=0.01$), JANET achieves better coverage under both predictors. Other methods again show significant under-coverage, particularly MC dropout.